OpenStudy (anonymous):

1+1=x x+x=? ?+?=?? ??+??=???

OpenStudy (anonymous):

too hard

OpenStudy (mimi_x3):

LOL, asking stupid questions like Harshir xDD

OpenStudy (anonymous):

ah that is so unusual only mimi can solve this one

OpenStudy (mimi_x3):

lmao, this time its too hard xD sorry i cant solve it :P:

OpenStudy (anonymous):

Word problem.

OpenStudy (anonymous):

2=x 2x=2(2) = 4 do this and u would get answer :) haha

OpenStudy (mimi_x3):

lols, there's no point of solving it for a stupid person :P:Pjkss

OpenStudy (anonymous):

U forgot...pleasssse heeelp!!!!

OpenStudy (anonymous):

very much hard only a grade 1st student can solve this

OpenStudy (anonymous):

pleasssse heeelp!!!!

OpenStudy (mimi_x3):

LOL, ishaan can help xDD

OpenStudy (anonymous):

And "See attached pic"

OpenStudy (anonymous):

kush did that unbelievable kush solved the question

OpenStudy (anonymous):

SEE ATTACHED PIC

OpenStudy (anonymous):

ah No i can't too tough for me

OpenStudy (anonymous):

How do I distribute the x?

OpenStudy (mimi_x3):

*thinking* Its so~ HARD to distribute x *sigh*

OpenStudy (anonymous):

only HARSHIR CAN SOLVE THIS!!!!!!

OpenStudy (mimi_x3):

YUP :D

OpenStudy (anonymous):

yea haha

OpenStudy (anonymous):

i cant solve it

OpenStudy (anonymous):

NOO HASHIR I KNOW U CAN BELIEVE!!!!!!

OpenStudy (anonymous):

gotta try for me bro

OpenStudy (mimi_x3):

LOL, his in kindergarten , its impossible for him to try

OpenStudy (anonymous):

yes !!!! poor kid

OpenStudy (anonymous):

or give me proof behind H +A=HA HA+S+H=HASH HASHIR=HARSHIR

OpenStudy (mimi_x3):

lols , HARSHIR is in an equation

OpenStudy (anonymous):

<$

OpenStudy (anonymous):

ok goota im sleepy

OpenStudy (anonymous):

night i mean day i mean w.e